• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name H. Ibrahim

    Log entry time 08:03:39 on October29,2002

    Entry number 89260

    keyword=Owl Shift Summary

    12:00 a.m. Owl shift takes over. Shift members: E. Voutier, P. Monaghan and H. Ibrahim.

    12:10 The momentum of the right arm for Q3D_f20r was not correct. We found that the regulator of the right arm dipole was turned off. Turning the regulator on solved the problem. Right arm momentum is now correct (2.617 GeV/c).

    12:45 a.m. We noticed that the beam current is not stable for the last couple of hours. We called MCC to know what the problem is and they told us that hall C requests are causing beam trips!

    02:00 a.m. Beam is better now.

    04:03 a.m. Target moved to Optics.

    04:31 a.m. Finished Q3D_f20r kinematics. Setting right arm momentum to 2.5897 GeV/c and angle to 39.51 to start next kinematics.

    04:40 a.m. Start Q3D_f30r kinematics by taking pointing data.

    05:00 a.m. Target moved to LD2 15 cm.

    08:00 a.m. End of owl shift.

    Continuing Q3D_f20r kinematics:

    Run # 2827 (Production, 1 M events, Beam was tripping all the time)
    Run # 2828 (Production, 1 M events, Beam quality was improved somewhat during this run)
    Run # 2829 (Production, 1 M events)
    Run # 2830 (Efficiency, Junk)
    Run # 2831 (Efficiency, 1 M events)
    Run # 2832 (Pointing, 315 k events)

    Q3D_f30r kinematics:

    Run # 2833 (Pointing, 352 k events)
    Run # 2834 (Efficiency, 1 M events)
    Run # 2835 (Production, 1 M events)
    Run # 2836 (Production, 1 M events)
    Run # 2837 (Production, 1 M events)
    Run # 2838 (Production, in progress)